Circular No. 7929

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2002dt IN ESO 516-G5 D. Hutchings and W. D. Li, University of California, report the discov-

ery by LOTOSS (cf. IAUC 7906) of an apparent supernova (mag ∼ 18.5) on unfiltered KAIT images taken on June 26.3 and 27.3 UT. SN 2002dt is located at α = 16^{h}03^{m}47^{s}.30, δ = −22^{o}44′19′′.5 (equinox 2000.0), which is 7′′.8 east and 11′′.1 south of the nucleus of ESO 516-G5. A KAIT image taken on June 17.3 showed nothing at this position (limiting mag ∼ 19.5). Li, R. Chornock, and A. V. Filippenko write that a spectrum of SN 2002dt, taken by M. M. Phillips with the Baade 6.5-m telescope at Las Campanas on June 28.0 UT, shows it to be clearly a supernova, but the exact type is uncertain. They add: “There are no hydrogen Balmer lines or the Si II 635.5-nm line that are typical of normal type-II and type-Ia supernovae, respectively. There are no Ca II H and K or infrared triplet lines that are typical of normal type-Ib/c supernovae. Narrow emission lines in the host galaxy yields a recession velocity of 7487 km/s. Converting to rest wavelengths, broad emission lines can be seen at 430, 503, and 560 nm. There are also dips around 580 and 700 nm. Based on the fact that the supernova is superposed on an H II region and lacks obvious type-Ia or type-II features, we suspect that this may be a core-collapse event related to the Ib/c variety (but possibly with substantial differences).”

SUPERNOVA 2002du G. Altavilla and colleagues (cf. IAUC 7780) report the discovery of a

supernova (V = 22.2) on CCD frames taken on Apr. 9.2 UT with the European Southern Observatory/Max-Planck-Institut 2.2-m telescope (+ Wide Field Imager). SN 2002du is located at α = 13^{h}53^{m}18^{s}.28, δ = −11^{o}37′28′′.8 (equinox 2000.0), which is within 1′′ of the nucleus of its host galaxy. Nothing was visible at this location on a similar image taken in Apr. 2001 (limiting mag ∼ 24). A spectrogram obtained on June 16.1 with the Very Large Telescope Yepun (+ FORS2) shows broad Hα with a P-Cyg profile, indicating that SN 2002du is a type-II supernova. Narrow emission lines of the parent galaxy provide a redshift of z = 0.21.

SUPERNOVA 2002ds IN ESO 581-G25 Li, Chornock, and Filippenko also report that a spectrum of SN 2002ds

(cf. IAUC 7928), taken as above by Phillips on June 28.0 UT, shows it to be a young type-II event, with strong hydrogen Balmer and He I 587.6-nm lines superposed on a blue continuum.

2002 June 29  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green